OpenStudy (anonymous):

Give reasons for each of the following steps: 35*100 = (3*10+5)100 = (3*10+5)10^2 = (3*10)10^2 + 5*10^2 = 3(10*10^2)+5*10^2 = 3*10^3 +5*10^2+ = 3*10^3+5*10^2+0*10+0*1 = 3500

OpenStudy (lopus):

35= 30 + 5 =\[3*10^{1} + 5*10^{0}\] because is decimal remenber that \[10^{0}=1\]

OpenStudy (lopus):

\[100=10^{2}=10*10\] (x+y)*z= (x*z)+(y*z)